### v2

@@ 4 Exact denoising of the RHM with message passing
In this section, we consider the denoising diffusion process for the RHM. Given a noisy observation $X^{(0)}=x(t)$ of the input variables at time $t$, we compute $p(x(0)|x(t))$ exactly, obtaining full control of the statistics of the ba diffusion process from time $t$ to time 00. In particular, given the tree structure of the model, we can compute the marginal probability of the values of all latent variables conditioned on $x(t)$ by using a message-passing algorithm. Therefore, we obtain the probability that a latent variable at level $\ell$ has changed when performing the forward-backward diffusion process for a duration $t$, a central quantity to interpret Figure 2.
For computing the marginal distributions, we use Belief Propagation (BP) Mossel (2001); Mezard & Montanari (2009), which gives exact results for a tree graph such as the Random Hierarchy Model. In this case, the leaves of the tree correspond to the input variables at the bottom layer, and the root corresponds to the class variable at the top of the hierarchy. Each rule connecting variables at different levels corresponds to a factor node, as shown in Figure 4.
@@
The input variables $X^{(0)}$, in their one-hot-encoding representation, undergo the forward diffusion process of Equation 2, which can be defined in continuous time and constant $\beta_{t}$ by redefining $\overline{\alpha_{t}}=e^{-2t}$ and taking the limit $T\rightarrow\infty$ Song et al. (2020).
The denoising is made in two steps: the initialisation of the messages at the leaves and the BP iteration.
Initialisation of the upward messages
In its one-hot-encoding representation, $X^{(0)}_{i}$ is a $v$-dimensional vector: taking the symbol $a_{\gamma}\in\{a_{1},\dots,a_{v}\}=\mathcal{A}$ corresponds to $X^{(0)}_{i}=e_{\gamma}$, with $e_{\gamma}$ a canonical basis vector. Its continuous diffusion process takes place in $\mathbb{R}^{v}$: given the value $X^{(0)}_{i}=x_{i}(t)$, we can compute the probability of its starting value $p(x_{i}(0)|x_{i}(t))$ using Bayes formula. As derived in Appendix A, we obtain
@@
with $\Delta_{t}=(1-\overline{\alpha_{t}})/\sqrt{\overline{\alpha_{t}}}$ and $Z=\sum_{\mu=1}^{v}e^{x_{i,\mu}(t)/\Delta_{t}}$. This computation is performed independently for each input variable $i$, and therefore does not take into account the spatial correlations given by the generative model. The probabilities of Equation 4 are used to initialise the BP upward messages ${\nu_{\uparrow}^{(0)}}=p(x_{i}(0)|x_{i}(t))$ at the input variables.
@@
The BP upward and downward iterations for the (unnormalised) upward and downward messages respectively read
@@
${\tilde{\nu}_{\downarrow}^{(\ell)}}(x_{1})$ $=\sum_{\begin{subarray}{c}x_{2},...,x_{s}\in\mathcal{A}^{\otimes(s-1)}\\ y\in\mathcal{A}\end{subarray}}{{\psi}^{(\ell+1)}}(y,x_{1},...,x_{s})$ (7)
@@
where $\nu_{\rho}^{(\ell)}(x)=\frac{\tilde{\nu}_{\rho}^{(\ell)}(x)}{\sum_{x^{\prime}}\tilde{\nu}_{\rho}^{(\ell)}(x^{\prime})}$, $\rho\in\{\uparrow,\downarrow\}$. The downward iteration, reported for $x_{1}$, can be trivially extended to the other variables $x_{i}$ by permuting the position indices. The values of ${\nu_{\uparrow}^{(0)}}(x_{i})$ and ${\nu_{\downarrow}^{(L)}}(y)$ are set by the initial conditions. In particular, we initialise ${\nu_{\uparrow}^{(0)}}(x_{i})$ as described in the previous paragraph and ${\nu_{\downarrow}^{(L)}}(y)=1/v$, which corresponds to a uniform prior over the possible classes $\mathcal{C}$ [^2] This assumption corresponds to unconditioned diffusion, where the DDPM is not biased towards any specific class.
Figure 5: Probability that the latent has not changed in the denoising process, corresponding to the largest marginal probability computed by BP, averaged for each layer, for varying inversion times of the diffusion process $t$. Data for the RHM with $v=32$, $m=8$, $s=2$, $L=10$. Each level of the tree, indicated in the legend, is represented with a different colour. We observe the same behaviour of the curves for ImageNet data in Figure 2: the probability of the correct class has a sharp transition at characteristic time scale, while the probabilities corresponding to latent variables in the lower levels change smoothly.
Exact solution We solve the BP upward and backward iterations numerically. In Figure 5, we show the probability corresponding to the correct symbol for each node of the tree. Remarkably, we note that (i) the probability for the correct class at layer $L$ displays a transition at a characteristic time which becomes sharper for increasing $L$, and (ii) the messages for the correct input variables and the correct latent variables at low levels of the tree change smoothly. In particular, the curves for messages at layer $L$ and layers $\ell<L$ invert their order at the transition, as in our observations on DDPMs and ImageNet data in Figure 2. This transition is our key finding, which we explain below.
@@ 5 Mean-field theory of denoising diffusion
In this section, we make a simplifying assumption for the initial noise acting on the input and adopt a mean-field approximation to justify the existence of a phase transition. Remarkably, this approximation turns out to be of excellent quality for describing the diffusion dynamics. Specifically, consider a reference configuration at the leaves variables $X_{i}^{(0)}={\overline{x}}_{i}$ that we would like to reconstruct, given a noisy observation of it. We assume that for each leaf variable, the noise is uniformly spread among the other symbols. [^3] This is a mild approximation, as documented in Appendix C In other words, our belief in the correct sequence is corrupted by $\epsilon\in[0,1]$:
@@
Hence, the initialisation condition of the upward BP messages at a leaf node $X_{i}^{(0)}$ becomes
@@
Figure 6: Phase diagram for inferring the class node using the upward iteration of BP. When $sf<1$, BP can infer the class if $\epsilon<\epsilon^{*}(sf)$. This transition is very well predicted by our theory. The inference region in the figure corresponds to the phase wherein the probability of the correct class is larger than the initialisation belief in the correct values of the leaves, that is $1-\epsilon+\frac{\epsilon}{v}$. Experimental data are for a single realisation of the RHM with $v=32$, $s=2$, $L=10$.
@@
We begin by studying the upward process from the leaves. Consider a true input tuple ${\overline{x}}_{1},\dots,{\overline{x}}_{s}$ which is associated with the higher-level feature ${\overline{y}}$. Given the randomness of the production rules, the messages are random variables depending on the specific realisation of the rules. We adopt a mean-field or annealed approximation that neglects the fluctuations coming from the random choice of rules. Specifically, we approximate the upward message by the average upward message exiting the corresponding factor node $\langle{\nu_{\uparrow}^{(1)}}(y)\rangle_{\psi}$ over the possible realisations of ${\psi}$. In Appendix B, we show that $\langle{\nu_{\uparrow}^{(1)}}(y)\rangle_{\psi}$ can take only two values: one for $y={\overline{y}}$ and one for $y\neq{\overline{y}}$, as expected by symmetry considerations. Therefore, mean messages have the same structure as Equation 10 and we can define a new $\epsilon^{\prime}$. Introducing the probability of reconstructions $p=1-\epsilon+\epsilon/v$ and $p^{\prime}=1-\epsilon^{\prime}+\epsilon^{\prime}/v$, we have
@@
Iterating this procedure across all the levels of the tree, we can compute the probability of recovering the correct class of the input. In particular, for large $L$, we are interested in studying the fixed points $p^{*}=F(p^{*})$ of the iteration map in Equation 11. As derived in Section B.1.1, when $sf>1$, this map has a repulsive fixed point $p^{*}=1$, which corresponds to $\epsilon=0$, and an attractive fixed point $p^{*}=\nicefrac{{1}}{{v}}$, corresponding to $\epsilon=1$. Thus, in this regime, inferring the class from the noisy observation of the input is impossible. In contrast, when $sf<1$, $p^{*}=1$ and $p^{*}=\nicefrac{{1}}{{v}}$ are both attractive fixed points, and a new repulsive fixed point $\nicefrac{{1}}{{v}}<p^{*}<1$ separating the other two emerges. Therefore, in this second regime, there is a phase transition between a phase in which the class can be recovered and a phase in which it cannot. These theoretical predictions are numerically confirmed in the phase diagram in Figure 6.
Physically, $sf<1$ corresponds to a regime in which errors at lower levels of the tree do not propagate: they can be corrected using information coming from neighbouring nodes, thanks to the fact that only a small fraction of the strings is consistent with the production rules of the generative model. Conversely, when $sf>1$, even small corruptions propagate through the entire tree up to the root node and BP cannot infer the class correctly.
@@
The same calculation can be repeated for the downward process, with the additional difficulty that the downward iteration mixes upward and downward messages. We refer the reader to Appendix B for the theoretical treatment.
@@
Combining the mean upward and downward messages, we obtain a theoretical prediction for the probabilities of reconstructing the correct values of the variables at each layer. We compare our theoretical predictions with numerical experiments in Figure 7. In these experiments, BP equations are solved exactly for a given RHM starting with the initialisation of Equation 10. Our theory perfectly captures the probability of reconstruction for the input nodes and the class. Moreover, Figure 14 in Appendix B shows that our theory predicts the probabilities of reconstruction of latent nodes at all layers.
@@
Similarly to our experiment on the ConvNeXt in Section 2, we investigate how the hidden representation of a model trained to classify the RHM changes when its input is denoised starting from a corruption noise $\epsilon$. We consider an instantiation of the RHM with $L=7$, $s=2$, $v=16$, and $m=4$. First, we train a convolutional neural network with $L=7$ layers, matching the tree structure of the model, with $n=300k$ training examples up to interpolation. The resulting architecture has 99.2% test accuracy. To sample new data from noisy observations of held-out data, we start by sampling the root using the marginal probability computed with BP. Then, we update the beliefs and the marginals conditioning on the sampled class, and sample one latent variable at layer $L-1$. We iterate this procedure node-by-node, descending the tree until we obtain a sampled configuration at the bottom layer Mezard & Montanari (2009). For each corrupting noise $\epsilon$ and each layer of the CNN, we compute the cosine similarity between post-activations for the initial and generated configurations. Figure 8 shows the obtained curves. Remarkably, we observe the same qualitative behaviour as in Figure 7, ultimately explaining the empirical observation of Figure 2.
Figure 7: Probability that the latent has not changed in the denoising process, corresponding to the largest marginal probability computed by BP, for varying $\epsilon$. Data for the RHM with $v=32$, $m=8$, $s=2$, $L=10$. Each level of the tree, indicated in the legend, is represented with a different colour. The black dashed lines are our mean-field theoretical predictions, which show excellent agreement with the experiments. In particular, the inversion between the curves for the top and bottom levels at the phase transition can be observed.
Figure 8: Cosine similarity between the post-activations following every layer of a deep CNN trained on the RHM ($v=16$, $m=4$, $s=2$, $L=7$) for the starting and sampled data. Each layer of the architecture, indicated in the legend, is represented with a different colour. The curves showcase the same inversion predicted by our theory (cf. Figure 7).
@@ 6 Conclusions
We have argued that reversing time in denoising diffusion models opens a window on the compositional nature of data. For synthetic hierarchical generative data models, low-level features can already change at small times, but the class remains most often the same. At larger times, a phase transition is found where the probability of remaining in the same class suddenly drops to random chance. Yet, low-level features identical to those of the initial sample can persist and compose the new sample. Strikingly, the same results are found for ImageNet. These observations may thus be more general, and testing them in text datasets, using diffusion language models He et al. (2022), is an interesting venue for future work.
@@
The interplay between the hierarchy in feature space and in time revealed here may help understand the puzzling success of diffusion models, including the number of data needed to train such methods, or why they can generalise and not simply memorise the empirical distribution on which they were trained Somepalli et al. (2022); Carlini et al. (2023); Yoon et al. (2023). More generally, our results put forward hierarchical generative models as tools to understand open questions for other methods, ranging from the emergence of new skills by the composition of more elementary ones in self-supervised learning to that of transferable representations in supervised or contrastive learning.
Impact statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Acknowledgements
We thank Francesco Cagnetta and Umberto Maria Tomasini for the helpful discussions. This work was supported by a grant from the Simons Foundation (#454953 Matthieu Wyart).
Appendix A Belief Propagation initialization for the denoising of the RHM
As discussed in Section 4, we define the diffusion process for the input variable $X_{i}^{(0)}$ in the space $\mathbb{R}^{v}$. In particular, its value $x(t)$ at time $t$ is
$x(t)=\sqrt{\overline{\alpha_{t}}}x(0)+\sqrt{1-\overline{\alpha_{t}}}\eta,$ (12)
with $\eta\sim\mathcal{N}(0,\mathbb{I}_{v})$ and $x(0)$ its starting value at time $t$, which is a one-hot-encoding vector of the form $x(0)=e_{\mu}$. Given the value $x(t)$, the conditional probabilities for the values of $x(0)$ are given by Bayes rule
$p\left(x(0)=e_{\mu}|x(t)\right)=\frac{p\left(x(t)|x(0)=e_{\mu}\right)\,p\left(x(0)=e_{\mu}\right)}{\sum_{\lambda}p\left(x(t)|x(0)=e_{\lambda}\right)\,p\left(x(0)=e_{\lambda}\right)}.$ (13)
The prior probabilities on $x(0)$ are taken to be uniform over the alphabet, i.e., $p\left(x(0)=e_{\lambda}\right)=1/v$, $\forall\lambda$, while $p\left(x(t)|x(0)=e_{\mu}\right)$ is given by the diffusion process of Eq. 12:
$\begin{split}p\left(x(t)|x(0)=e_{\mu}\right)&=C_{t}\exp\left[-\frac{1}{2(1-\overline{\alpha_{t}})}\sum_{\gamma}\left(x_{\gamma}(t)-\sqrt{\overline{\alpha_{t}}}e_{\mu}\right)^{2}\right]=\\ &=C_{t}\exp\left[-\frac{\left(\|x(t)\|^{2}-\overline{\alpha_{t}}\right)}{2(1-\overline{\alpha_{t}})}\right]\exp\left[\frac{\sqrt{\overline{\alpha_{t}}}}{1-\overline{\alpha_{t}}}x_{\mu}(t)\right],\end{split}$ (14)
where $C_{t}$ is the normalization constant. Putting Eq. 14 into Eq. 13, we obtain
$p\left(x(0)=e_{\mu}|x(t)\right)=\frac{1}{Z}e^{\frac{\sqrt{\overline{\alpha_{t}}}}{1-\overline{\alpha_{t}}}x_{\mu}(t)},$ (15)
with $Z=\sum_{\lambda=1}^{v}e^{\frac{\sqrt{\overline{\alpha_{t}}}}{1-\overline{\alpha_{t}}}x_{\lambda}(t)}$.
Appendix B Belief Propagation equations
Given a factor tree-graph, the Belief Propagation (BP) equations compute iteratively the messages going from the variable nodes to the factor nodes and vice-versa, starting from the initialization conditions at the leaves and root of the tree-graph Mezard & Montanari (2009). For the generative model defined in Sec. 3, the leaves correspond to the variables at the bottom layer while the root is the class variable at the top of the hierarchy. Each rule, connecting variables at different layers, corresponds to a factor node. The BP messages that flow from the variable nodes to the factor nodes, therefore, correspond to upward messages, while those going from factor nodes to variables correspond to downward messages (Fig. 9).
To each variable node $X^{(\ell)}_{i}$ at level $\ell$, we associate the upward messages ${\nu_{\uparrow}^{(\ell)}}$ and downward messages ${\nu_{\downarrow}^{(l)}}$, one for each possible value of the alphabet it can take. To simplify the notation, here we consider how messages propagate from one level to the other and we call $Y$ the variable corresponding to the higher level and $X_{i=1,...,s}$ the lower level ones connected to it. The factor node connecting them is such that, for each possible association $y\rightarrow x_{1},\dots,x_{s}$, it takes values
${{\psi}^{(\ell)}}(y,x_{1},...,x_{s})=\begin{cases}1,\quad\text{if }y\rightarrow(x_{1},...,x_{s})\text{ is a rule at layer }\ell\\ 0,\quad\text{otherwise}.\end{cases}$
The BP upward and downward iterations are defined as follows.
  • Upward iteration:
    ${\tilde{\nu}_{\uparrow}^{(\ell+1)}}(y)=\sum_{x_{1},...,x_{s}\in\mathcal{A}^{\otimes s}}{{\psi}^{(\ell+1)}}(y,x_{1},...,x_{s})\prod_{i=1}^{s}{\nu_{\uparrow}^{(\ell)}}(x_{i})\ ,$ (16)
    ${\nu_{\uparrow}^{(\ell)}}(y)=\frac{{\tilde{\nu}_{\uparrow}^{(\ell)}}(y)}{\sum_{y^{\prime}}{\tilde{\nu}_{\uparrow}^{(\ell)}}(y^{\prime})}.$ (17)
  • Downward iteration:
    ${\tilde{\nu}_{\downarrow}^{(\ell)}}(x_{1})=\sum_{\begin{subarray}{c}x_{2},...,x_{s}\in\mathcal{A}^{\otimes(s-1)}\\ y\in\mathcal{A}\end{subarray}}{{\psi}^{(\ell+1)}}(y,x_{1},...,x_{s})\ {\nu_{\downarrow}^{(\ell+1)}}(y)\prod_{i=2}^{s}{\nu_{\uparrow}^{(\ell)}}(x_{i})$ (18)
    ${\nu_{\downarrow}^{(l)}}(x)=\frac{{\tilde{\nu}_{\downarrow}^{(\ell)}}(x)}{\sum_{x^{\prime}}{\tilde{\nu}_{\downarrow}^{(\ell)}}(x^{\prime})}.$ (19)
${\nu_{\uparrow}^{(\ell)}}(y)$ and ${\nu_{\downarrow}^{(\ell)}}(x)$ are fluctuating quantities that depend on the position of the node.

### v3

@@ 3 Optimal denoising of the RHM with message passing
In this section, we characterize the Bayes optimal denoising process for the RHM. Given a noisy observation $X^{(0)}=x(t)$ of the input variables at time $t$, we compute $p(x(0)|x(t))$ exactly, obtaining full control of the statistics of the backward diffusion process from time $t$ to time 00. In particular, given the tree structure of the model, we can compute the marginal probability of the values of all latent variables conditioned on $x(t)$ by using a message-passing algorithm. Therefore, we obtain the probability that a latent variable at level $\ell$ has changed when performing the forward-backward diffusion process for a duration $t$, a central quantity to interpret Figure 2. The optimal denoising corresponds to reconstructing the data distribution $p(x(0))$ exactly. This perfect reconstruction corresponds to a diffusion model achieving perfect generalization. Although this is a strong assumption for modeling a neural network trained with empirical risk minimization, like the one considered in Section 1, our theoretical analysis captures the phenomenology of our experiments.
Belief Propagation
For computing the marginal distributions, we use Belief Propagation (BP) (52, 53), which gives exact results for a tree graph such as the Random Hierarchy Model. In this case, the leaves of the tree correspond to the input variables at the bottom layer, and the root corresponds to the class variable at the top of the hierarchy. Each rule connecting variables at different levels corresponds to a factor node, as shown in Figure 4.
@@
The input variables $X^{(0)}$, in their one-hot-encoding representation, undergo the forward diffusion process of Equation 2, which can be defined in continuous time and constant $\beta_{t}$ by redefining $\overline{\alpha_{t}}=e^{-2t}$ and taking the limit $T\rightarrow\infty$ (14).
The denoising is made in two steps: the initialization of the messages at the leaves and the BP iteration.
initialization of the upward messages
In its one-hot-encoding representation, $X^{(0)}_{i}$ is a $v$-dimensional vector: taking the symbol $a_{\gamma}\in\{a_{1},\dots,a_{v}\}=\mathcal{A}$ corresponds to $X^{(0)}_{i}=e_{\gamma}$, with $e_{\gamma}$ a canonical basis vector. Its continuous diffusion process takes place in $\mathbb{R}^{v}$: given the value $X^{(0)}_{i}=x_{i}(t)$, we can compute the probability of its starting value $p(x_{i}(0)|x_{i}(t))$ using Bayes formula. As derived in SI Appendix 1, we obtain
@@
with $\Delta_{t}=(1-\overline{\alpha_{t}})/\sqrt{\overline{\alpha_{t}}}$ and $Z=\sum_{\mu=1}^{v}e^{x_{i,\mu}(t)/\Delta_{t}}$. This computation is performed independently for each input variable $i$, and therefore does not take into account the spatial correlations given by the generative model. The probabilities of Equation 4 are used to initialize the BP upward messages ${\nu_{\uparrow}^{(0)}}=p(x_{i}(0)|x_{i}(t))$ at the input variables.
@@
The BP upward and downward iterations for the (unnormalized) upward and downward messages respectively read
@@
${\tilde{\nu}_{\downarrow}^{(\ell)}}(x_{1})$ $=\sum_{\begin{subarray}{c}x_{2},...,x_{s}\in\mathcal{A}^{\otimes(s-1)}\\ y\in\mathcal{A}\end{subarray}}{{\psi}^{(\ell+1)}}(y,x_{1},...,x_{s})$
@@
where $\nu_{\rho}^{(\ell)}(x)=\frac{\tilde{\nu}_{\rho}^{(\ell)}(x)}{\sum_{x^{\prime}}\tilde{\nu}_{\rho}^{(\ell)}(x^{\prime})}$, $\rho\in\{\uparrow,\downarrow\}$. The downward iteration, reported for $x_{1}$, can be trivially extended to the other variables $x_{i}$ by permuting the position indices. The values of ${\nu_{\uparrow}^{(0)}}(x_{i})$ and ${\nu_{\downarrow}^{(L)}}(y)$ are set by the initial conditions. In particular, we initialize ${\nu_{\uparrow}^{(0)}}(x_{i})$ as described in the previous paragraph and ${\nu_{\downarrow}^{(L)}}(y)=1/v$, which corresponds to a uniform prior over the possible classes $\mathcal{C}$ [^2] This assumption corresponds to unconditioned diffusion, where the DDPM is not biased towards any specific class.
Figure 5: Probability that the latent has not changed in the denoising process, corresponding to the largest marginal probability computed by BP, averaged for each layer, for varying inversion times of the diffusion process $t$. Data for the RHM with $v=32$, $m=8$, $s=2$, $L=10$. Each level of the tree, indicated in the legend, is represented with a different color. We observe the same behavior of the curves for ImageNet data in Figure 2: the probability of the correct class has a sharp transition at a characteristic time scale, while the probabilities corresponding to latent variables in the lower levels change smoothly.
Results
We run the BP upward and backward iterations numerically. In Figure 5, we show the probability corresponding to the correct symbol for each node of the tree. Remarkably, we note that (i) the probability for the correct class at layer $L$ displays a transition at a characteristic time which becomes sharper for increasing $L$, and (ii) the messages for the correct input variables and the correct latent variables at low levels of the tree change smoothly. In particular, the curves for messages at layer $L$ and layers $\ell<L$ invert their order at the transition, as in our observations on DDPMs and ImageNet data in Figure 2. This transition is one of our key findings, which we explain below.
@@ 4 Mean-field theory of denoising diffusion
In this section, we make a simplifying assumption for the initial noise acting on the input and adopt a mean-field approximation to justify the existence of a phase transition. Remarkably, this approximation turns out to be of excellent quality for describing the diffusion dynamics. Specifically, consider a reference configuration at the leaves variables $X_{i}^{(0)}={\overline{x}}_{i}$ that we would like to reconstruct, given a noisy observation of it. We assume that for each leaf variable, the noise is uniformly spread among the other symbols. [^3] This is a mild approximation, as documented in SI Appendix 3. In other words, our belief in the correct sequence is corrupted by $\epsilon\in[0,1]$:
@@
Hence, the initialization condition of the upward BP messages at a leaf node $X_{i}^{(0)}$ becomes
@@
Figure 6: Phase diagram for inferring the class node using the upward iteration of BP. When $sf<1$, BP can infer the class if $\epsilon<\epsilon^{*}(sf)$. This transition is very well predicted by our theory. The inference region in the figure corresponds to the phase wherein the probability of the correct class is larger than the initialization belief in the correct values of the leaves, that is $1-\epsilon+\frac{\epsilon}{v}$. Experimental data are for a single realization of the RHM with $v=32$, $s=2$, $L=10$.
@@
We begin by studying the upward process from the leaves. Consider a true input tuple ${\overline{x}}_{1},\dots,{\overline{x}}_{s}$ which is associated with the higher-level feature ${\overline{y}}$. Given the randomness of the production rules, the messages are random variables depending on the specific realization of the rules. We adopt a mean-field or annealed approximation that neglects the fluctuations coming from the random choice of rules. Specifically, we approximate the upward message by the average upward message exiting the corresponding factor node $\langle{\nu_{\uparrow}^{(1)}}(y)\rangle_{\psi}$ over the possible realizations of ${\psi}$. In SI Appendix 2, we show that $\langle{\nu_{\uparrow}^{(1)}}(y)\rangle_{\psi}$ can take only two values: one for $y={\overline{y}}$ and one for $y\neq{\overline{y}}$, as expected by symmetry considerations. Therefore, mean messages have the same structure as Equation 7 and we can define a new $\epsilon^{\prime}$. Introducing the probability of reconstructions $p=1-\epsilon+\epsilon/v$ and $p^{\prime}=1-\epsilon^{\prime}+\epsilon^{\prime}/v$, we have
@@
Iterating this procedure across all the levels of the tree, we can compute the probability of recovering the correct class of the input. In particular, for large $L$, we are interested in studying the fixed points $p^{*}=F(p^{*})$ of the iteration map in Equation 8. As derived in SI Appendix 2.A, when $sf>1$, this map has a repulsive fixed point $p^{*}=1$, which corresponds to $\epsilon=0$, and an attractive fixed point $p^{*}=\nicefrac{{1}}{{v}}$, corresponding to $\epsilon=1$. Thus, in this regime, inferring the class from the noisy observation of the input is impossible. In contrast, when $sf<1$, $p^{*}=1$ and $p^{*}=\nicefrac{{1}}{{v}}$ are both attractive fixed points, and a new repulsive fixed point $\nicefrac{{1}}{{v}}<p^{*}<1$ separating the other two emerges. Therefore, in this second regime, there is a phase transition between a phase in which the class can be recovered and a phase in which it cannot. These theoretical predictions are numerically confirmed in the phase diagram in Figure 6.
Physically, $sf<1$ corresponds to a regime in which errors at lower levels of the tree do not propagate: they can be corrected using information coming from neighboring nodes, thanks to the fact that only a small fraction of the strings are consistent with the production rules of the generative model. Conversely, when $sf>1$, even small corruptions propagate through the entire tree up to the root node and BP cannot infer the class correctly.
@@
The same calculation can be repeated for the downward process, with the additional difficulty that the downward iteration mixes upward and downward messages. We refer the reader to SI Appendix 2 for the theoretical treatment.
@@
Combining the mean upward and downward messages, we obtain a theoretical prediction for the probabilities of reconstructing the correct values of the variables at each layer. We compare our theoretical predictions with numerical experiments in Figure 7-(a). In these experiments, BP equations are solved exactly for a given RHM starting with the initialization of Equation 7. Our theory perfectly captures the probability of reconstruction for the input nodes and the class. Moreover, in SI Appendix 2 we show that our theory predicts the probabilities of reconstruction of latent nodes at all layers.
@@
Similarly to our experiment on the ConvNeXt in Section 1, we investigate how the hidden representation of a model trained to classify the RHM changes when its input is denoised starting from a corruption noise $\epsilon$. We consider an instantiation of the RHM with $L=7$, $s=2$, $v=16$, and $m=4$. First, we train a convolutional neural network with $L=7$ layers, matching the tree structure of the model, with $n=300k$ training examples up to interpolation. The resulting architecture has 99.2% test accuracy. To sample new data from noisy observations of held-out data, we start by sampling the root using the marginal probability computed with BP. Then, we update the beliefs and the marginals conditioning on the sampled class, and sample one latent variable at layer $L-1$. We iterate this procedure node-by-node, descending the tree until we obtain a sampled configuration at the bottom layer (53). For each corrupting noise $\epsilon$ and each layer of the CNN, we compute the cosine similarity between post-activations for the initial and generated configurations. Panel (b) of Figure 7 shows the obtained curves. Remarkably, we observe the same qualitative behavior as in panel (a) of Figure 7, ultimately explaining the empirical observation of Figure 2.
Figure 7: (a) Probability that the latent has not changed in the denoising process, corresponding to the largest marginal probability computed by BP, for varying $\epsilon$. Data for the RHM with $v=32$, $m=8$, $s=2$, $L=10$. Each level of the tree, indicated in the legend, is represented with a different color. The black dashed lines are our mean-field theoretical predictions, which show excellent agreement with the experiments. In particular, the inversion between the curves for the top and bottom levels at the phase transition can be observed. (b) Cosine similarity between the post-activations following every layer of a deep CNN trained on the RHM ($v=16$, $m=4$, $s=2$, $L=7$) for the starting and sampled data. Each layer of the architecture, indicated in the legend, is represented with a different color. The curves showcase the same inversion predicted by our theory (cf. panel (a)).
@@ 5 Conclusions
We have argued that reversing time in denoising diffusion models opens a window on the compositional nature of data. For synthetic hierarchical generative models of data, where the Bayes optimal denoising can be exactly computed, low-level features can already change at small times, but the class remains most often the same. At larger times, a phase transition is found where the probability of remaining in the same class suddenly drops to random chance. Yet, low-level features identical to those of the initial sample can persist and compose the new sample. Strikingly, this theoretical analysis characterizes well the results found with ImageNet, where the denoising is performed by a trained U-Net. Interestingly, the structure of the U-Net with the skip connections between the encoder and decoder parts mimics the upward and downward iterations of belief propagation, where the downward process mixes upward and downward messages. In fact, building on the present work (54) shows that U-Nets are capable of effectively approximating the belief propagation denoising algorithm. Investigating whether the function learned by U-Nets approximates BP is a promising avenue for future work. In the present work, we used the internal representation of deep networks as a proxy for the hierarchical structure of images. An interesting direction for future work will be using deep hierarchical segmentation techniques (55, 56, 57, 58) to extract latent variables, so as to test our predictions on their evolution in forward-backward experiments. Finally, future work can test our theoretical predictions on other modalities successfully handled by diffusion models, such as language and biological structures.
@@
The interplay between the hierarchy in feature space and in time revealed here may help understand the puzzling success of diffusion models, including the number of data needed to train such methods, or why they can generalize and not simply memorize the empirical distribution on which they were trained (59, 60, 61). More generally, our results put forward hierarchical generative models as tools to understand open questions for other methods, ranging from the emergence of new skills by the composition of more elementary ones in foundation models to that of transferable representations in self-supervised learning.
[…]
The code for reproducing the experiments is available at github.com/pcsl-epfl/forward-backward-diffusion (42).